D-Time AO

Software changes tonight:

WFS

From old to new version:

log into soul-sxwfs

w_stop_all

cd /home/AOeng/SOUL.latest

source flao_environment.sh

sudo su

cp /etc/init/keepUp.conf.new /etc/init/keepUp.conf

exit

wfs_init

# all the GUIs must be then launched from this terminal or another where you did:

cd /home/AOeng/SOUL.latest

source flao_environment.sh

From new to old version:

log into soul-sxwfs

cd /home/AOeng/SOUL.latest

source flao_environment.sh

w_stop_all

sudo su

cp /etc/init/keepUp.conf.old /etc/init/keepUp.conf

exit

cd /home/AOeng/SOUL

source flao_environment.sh

wfs_init

AdSec

Instructions to activate the "flao" environment (note: please keep the wind on the AdSec low during those steps):
  1. uncomment and comment the lines in the ~/.bashrc file as below.
  2. Power off the unit, and adsc_stop (note: the adsec will be on TSS, please do the following steps as soon as possible)
  3. kill all the sx adsec terminals
  4. login again
  5. adsc_start and power on

uncomment:
if [ -f ~/soul/soul.source/flao_environment.sh ]; then
. ~/soul/soul.source/flao_environment.sh
fi
comment:
#if [ -f ~/soul/flao_environment.sh ]; then
# . ~/soul/flao_environment.sh
#fi

Had to do a quick Kludge with .bashrc, once we have just one environment, we should't have this issue anymore.

[AOeng@soul-sxwfs 29]$ cat ~/.bashrc
# .bashrc

# Source global definitions
if [ -f /etc/bashrc ]; then
. /etc/bashrc
fi

unalias ls
alias ls="ls -F"
alias go2data="cd /local/aomeas/adsec_data/"
alias go2calib="cd /local/aomeas/adsec_calib/"
alias go2aotab="cd /home/AOeng/aoroot/calib/wfs/current/ao/"
alias go2aoconf="cd /home/AOeng/aoroot/conf/wfs/current/processConf/"

# User specific aliases and functions
#if [ -f ~/SOUL/flao_environment.sh ]; then
# . ~/SOUL/flao_environment.sh
#fi

# User specific aliases and functions
if [ -f ~/SOUL.latest/flao_environment.sh ]; then
. ~/SOUL.latest/flao_environment.sh
source ~/SOUL/flao_environment.sh
source ~/SOUL.latest/flao_environment.sh
fi

08:23
  • WFS started up after the first 3 tries.
08:26
  • OSCO #13
  • Had to shift the pupils for bin1 back to the original values.
  • The pupil is still 1 pixel above the where we had it in February.
  • We're going to have to make new pupil files because I don't have any that will work.
  • We're going to do ESM later, but we'll be using bin4, which should be OK.
09:00
  • Fabio is going to make some changes to auxloops and optical loop diagnostic, to get seeing that is corrected for elevation and something else that I didn't hear.
  • Interesting that we got one skip frame, basically, every couple of min. The seeing is good, but variable, so it could be that the changing of the GOPT gain caused a quick safe skip? Not a problem, just weird.
  • housekeeper.L |INF| 630|2021-03-29 07:59:13.821423| SwitchSafeSkipFrameCnt > 0
    housekeeper.L |INF| 35012|2021-03-29 08:56:25.900053| SwitchSafeSkipFrameCnt > 1
    housekeeper.L |INF| 35100|2021-03-29 08:56:34.502296| SwitchSafeSkipFrameCnt > 2
    housekeeper.L |INF| 38704|2021-03-29 09:02:35.580558| SwitchSafeSkipFrameCnt > 3
    housekeeper.L |INF| 39614|2021-03-29 09:04:07.568806| SwitchSafeSkipFrameCnt > 4
  • perl -ne '$bob = $ENV{bob}; if(/SafeSkipFrameCnt\s\>\s(.+)$/){if($1 = $bob){print $_; $bob = $1; $ENV{bob} = $bob}}'
09:07
  • Going to do ESM.
  • But first a standard star.
09:30
  • Heading to the dwarf galaxy.........
09:38
  • CCD 47 timed out.
  • This seems to have been caused by the BCU47 being down, as seen in the HW GUI.
  • I reset the bcu47 in the hardware GUI, under the power controller, and that seems to have done the trick.
09:41
  • Bin 4, 359 Hz
  • Held the gain, so if the loop opens....
  • I had to manually set the TT gain because the optimize gain process set it to 0.05, which is absurd.
  • The minimum on the graph was more like 0.5, so I set it to 0.35.
  • Gains, 0.35 0.20 0.10 (ho2 doesn't count)
  • Seeing holding at about 1.2
  • 20210329_094503, 20210329_094549, 20210329_094737, 20210329_094800, 20210329_094825, 20210329_094914
  • 20210329_095633, 20210329_100025, 20210329_100052, 20210329_100113, 20210329_100242, 20210329_100323
10:50
  • Lost the loop while we were transiting. We'll re-close shortly.
10:56
  • Back in action!!!
  • We ran into the Rerotator limit. Rotating the LUCI 1 rotator by 360 seems to have taken care of that.
11:54
  • Implementing Fabio's software change, we stopped and started the WFS software so many times that the adsec processes disconnected.
  • Master Diagnostic is stuck in "Init".
  • Alfio is working on it. Possibly the software change had something to do with it.
  • The process could not create a shared buffer.
  • Alfio go the shared buffer to work by restarting as the old system then restarting as the new system.
  • 20210329_124221

-- GregoryTaylor - 29 Mar 2021
Topic revision: r1 - 29 Mar 2021, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback